Skip to content
CalcGospel 國際數學圖譜
返回

CIE 9709 2026 June Paper 52 Q5

A Level / CIE / S1

CIE 9709 2026 June Paper 52 Paper · Question 5

题目

Problem

A box contains 3 red pens, 5 blue pens and 2 green pens. Lui chooses one pen at random from the box and keeps it.

A green pen is added to the box. Mira now chooses one pen at random from the box and keeps it.

(a) Find the probability that Lui and Mira both have a blue pen.

(1)

(b) Find the probability that Mira’s pen is red.

(3)

(c) Find the probability that Lui’s pen is red given that Mira’s pen is blue or green.

(3)
题目中文翻译

一个盒子里有 3 支红笔、5 支蓝笔和 2 支绿笔。Lui 从盒中随机取出一支笔并保留。

随后向盒中加入一支绿笔。Mira 再从盒中随机取出一支笔并保留。

(a) 求 Lui 和 Mira 取到的都是蓝笔的概率。

(1)

(b) 求 Mira 取到红笔的概率。

(3)

(c) 已知 Mira 取到蓝笔或绿笔,求 Lui 取到红笔的概率。

(3)
原卷通用作答要求

Answer all questions. You should use a calculator where appropriate.

You must show all necessary working clearly; no marks will be given for unsupported answers from a calculator.

Give non-exact numerical answers correct to 3 significant figures, or 1 decimal place for angles in degrees, unless a different level of accuracy is specified in the question.

请回答所有题目。适当时应使用计算器。必须清楚展示所有必要的计算过程;仅给出计算器答案而没有过程,不予评分。除非题目另有精度要求,非精确数值答案保留 3 位有效数字,以度为单位的角度保留 1 位小数。

解答

(a)

解法一

思路

展开

先取走一支蓝笔,加入的是绿笔,因此第二次抽取时仍有10支笔,但蓝笔只剩4支。

答题过程

展开 P(both blue)=510×410=15.\mathrm{P}(\text{both blue})=\frac5{10}\times\frac4{10} =\boxed{\frac15}.

(b)

解法一

思路

展开

按Lui取到的颜色分类。每一种情况下先更新红笔数量;取走一支再加入一支,第二次的总数仍为10。

答题过程

展开

Write R1,B1,G1R_1,B_1,G_1 for Lui’s colour and R2R_2 for Mira obtaining red.

If Lui takes red, 2 red pens remain; otherwise, 3 remain. Adding a green pen restores the total number of pens to 10. Hence

P(R2)= P(R1)P(R2∣R1) +P(B1)P(R2∣B1) +P(G1)P(R2∣G1)= 310210+510310+210310= 27100.\begin{align*} \mathrm{P}(R_2)=&\,\mathrm{P}(R_1)\mathrm{P}(R_2\mid R_1)\\[4mm] &\,\hspace{2pt}+\mathrm{P}(B_1)\mathrm{P}(R_2\mid B_1)\\[4mm] &\,\hspace{4pt}+\mathrm{P}(G_1)\mathrm{P}(R_2\mid G_1)\\[4mm] =&\,\frac3{10}\frac2{10}+\frac5{10}\frac3{10} +\frac2{10}\frac3{10}\\[4mm] =&\,\boxed{\frac{27}{100}}. \end{align*}

(c)

解法一

思路

展开

条件是Mira取蓝或绿,分母可由(b)取补集。分子是Lui取红且Mira不取红,需要按取红后更新的笔数计算。

答题过程

展开

If Lui takes red, Mira has 8 non-red pens out of 10 to choose from. Therefore,

P(R1∩(B2∪G2))=310×810=24100.\mathrm{P}\bigl(R_1\cap(B_2\cup G_2)\bigr) =\frac3{10}\times\frac8{10}=\frac{24}{100}.

From (b),

P(B2∪G2)=1−27100=73100.\mathrm{P}(B_2\cup G_2)=1-\frac{27}{100}=\frac{73}{100}.

Thus

P(R1∣B2∪G2)=24/10073/100=2473.\mathrm{P}(R_1\mid B_2\cup G_2) =\frac{24/100}{73/100} =\boxed{\frac{24}{73}}.